OpenStudy (boldgeorge):

Medal for anyone who can help! Three bricklayers and four carpenters earn a total of $910 per day. Ten bricklayers and five carpenters earn a total of $1850 a day. a) Set up a system of equations to describe the word problem. b) Is there an advantage to eliminating either variable first? c) Determine the daily wage of a bricklayer.

OpenStudy (anonymous):

let b equal the daily salary of one bricklayer and c for carpenters 3b + 4c = 910 10b + 5c = 1850 the advantage of eliminating c first is that we'll be left with b and we can answer the question of how many bricklayers; they didn't ask about carpenters. also c is easiest to work with since in equation 2 you can isolate it by dividing it entirely by 5. re-written as: 3b + 4c = 910 2b + c = 370 substituion: second line: c = 370 - 2b plug that in the first line: 3b + 4(370 - 2b) = 910 solve for b :)

OpenStudy (boldgeorge):

b=114?

OpenStudy (anonymous):

yup :)

OpenStudy (boldgeorge):

sweet! thanks for the help!